# TAMO:Fine-Grained Root Cause Analysis via Tool-Assisted LLM Agent with Multi-Modality Observation Data

### v2

@@ Abstract
With the development of distributed systems, microservices and cloud native technologies have become central to modern enterprise software development. Despite bringing significant advantages, these technologies also increase system complexity and operational challenges. Implementing automated root cause analysis to reduce manual intervention has become a key aspect of modern software operations and maintenance. In recent years, large language models (LLMs) have made breakthroughs in root cause context inference and the integration of domain-specific knowledge, providing new solutions for Artificial Intelligence for IT Operations (AIOps). However, existing LLM-based approaches face three key challenges: context window limitation, modal input constraint, and dynamic dependence hallucination. To address these issues, we propose a tool-assisted LLM agent with multi-modality observation data, namely TAMO, for fine-grained RCA. It unifies multi-modal observational data into time-aligned representations to extract consistent features and employs specialized root cause localization and fault classification tools for perceiving the contextual environment. This approach overcomes the limitations of LLM in handling real-time changing service dependencies and raw observational data and guides LLM to generate repair strategies aligned with system contexts by structuring key information into a prompt. Experimental results show that TAMO performs well in root cause analysis when dealing with public datasets characterized by heterogeneity and common fault types, demonstrating its effectiveness.
Index Terms:
Root cause analysis, Tool-Assisted LLM agent, Cloud-native systems, Multimodal data, Diffusion.
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ I Introduction
In recent years, Microservice architecture and cloud-native technologies have become foundational elements in modern enterprise software development, driving the rapid development of distributed systems. [28],[32],[39]. Microservices decompose traditional monolithic applications into independent distributed services, while cloud native technologies leverage containerization and container orchestration[21] to enhance the capabilities for flexible resource allocation, dynamic scaling, and continuous delivery[30]. In enterprise production environments, microservice systems are composed of numerous independent services and containerized entities[45], which significantly increased system complexity, These systems have intricate inter-service dependencies and communication patterns, where the failure of any component (e.g., network resource corruption) can trigger cascading effects that jeopardize overall system stability, potentially leading to significant economic losses[2]. How to achieve accurate root cause identification and resolution in complex distributed systems has become a key factor in reducing production risks and minimizing losses for modern enterprises.
@@
Root cause analysis aims to quickly correlate related multiple entity features, identify the root cause and determine the corresponding fault type, and provide actionable solutions, thus realizing an automatic fault response mechanism and minimizing human intervention[42]. How to perform an accurate and efficient root cause analysis in the distributed environment has become an important research direction. Traditional RCA methods typically rely on rule-based anomaly detection or statistical correlation analysis (e.g. causal graphs[24]) to localize faults. These methods struggle to jointly analyze heterogeneous data modalities (e.g., unstructured logs and structured metrics), often leading to fragmented information and frequent error detection. DL-based RCA methods attempt to incorporate multimodal feature fusion and dependency-sensitive Graph Neural Networks (GNNs)[23][44] to achieve accurate fault localization. Although these methods demonstrate good performance in specific tasks such as fault localization in complex systems, they still face significant challenges in real-time analysis and fault interpretability, heavily relying on the expertise of engineers to complete the root cause analysis and remediation process. Recently, large language models have demonstrated revolutionary capabilities in automated root cause analysis, opening new avenues for AIOps [9], [12]. LLM can combine system faults with the current system context to generate actionable fault diagnosis reports, as evidenced by prototypes such as RCACopilot[6]. Recent works such as RCAgent[35] and mABC[43] developed LLM-driven automated RCA frameworks that enable accurate fault localization and provide reasonable repair suggestions, thus further advancing the field of automated operations and maintenance.
However, although existing LLM-based methods have significantly alleviated the reliance on expert knowledge in traditional DL-based methods, they still face the following three key challenges in real distributed system environments:
  • Context Window Limitation: Limited by fixed-length context windows, the LLM is unable to effectively process massive amounts of contextual monitoring data in real-time. This constraint often leads to data truncation or compression during analysis, making it difficult to utilize historical information to detect anomalous trends in system components, thereby affecting the rationale and accuracy of root cause analysis. So how to extract useful information from the complex system context and input it into the LLM with limited context length?
  • Modal Input Constraint: Existing large language models such as ChatGPT[1], DeepSeek[27], Qwen[3], etc., either only support text data as input or enable modality input by mapping different modal features to a unified textual semantic space. However, these models often suffer from significant key feature loss when dealing with observational data, especially metric data, due to the lack of a obvious correlation between time series data and semantic text, which hinders the ability to analyze and reason effectively. So how to efficiently unify multiple modalities of observation data to assist the LLM in providing accurate root cause analysis results?
  • Dynamic Dependence Hallucination: The LLM is based on static, pre-trained knowledge, which makes it challenging to model service dependencies in systems. It can only simulate simple graph structures using text prompts. As noted in [34], LLMs struggle with reasoning over complex graphs, which limits their ability to effectively capture service dependencies. Additionally, the dynamic nature of service call chains in the system means that evolving dependencies can cause the LLM to generate hallucination that contradict actual dependencies, leading to inaccurate root cause analysis. So how to assist the LLM in modeling fault propagation paths using call chain relationships to achieve fine-grained root cause localization?
@@
To address these limitations, we propose a tool-assisted LLM agent framework based on domain-specific perceptive models. This framework decouples the LLM from the raw observational data by utilizing specialized observational data analysis tools($\mathscr{T}_{1}$-$\mathscr{T}_{3}$), addressing the constraints of text input and the limitations of context in the LLM. These specialized model tools are used to perceive the contextual environment, and their results form structured text inputs that are fed back to the RCA expert agent for further analysis, generating accurate root cause analysis results and reasonable maintenance recommendations. Specifically, to address the problem of multimodal data alignment, we propose a dual-branch diffusion model as the data feature observation tool ($\mathscr{T}_{1}$), which unifies multimodal observability data (logs, traces, metrics) into temporally consistent representations. To address the issue of dynamic dependencies, the root cause localization tool ($\mathscr{T}_{2}$) performs a joint analysis of entity features and call topologies, modeling the fault propagation chain, and identifying the root cause through frequency-domain attention-driven temporal causal analysis. Subsequently, the fault classification tool ($\mathscr{T}_{3}$) further analyzes the fault and maps the symptoms to predefined fault types. The main contributions are summarized as follows:
  • We propose a tool-assisted LLM agent framework, which integrates domain-specific model tools with LLMs for root cause analysis of microservices. This framework enables the LLM to analyze the system state in real time by calling perception tools, decoupling the agent from the raw observational data, allowing it to generate accurate root cause analysis results and reasonable maintenance recommendations.
  • We present a dual-branch collaborative diffusion tool that captures potentially consistent patterns across modalities and provides a unified time series representation for downstream analysis tasks through condition-oriented collaborative reconstruction.
@@
  • We design a root cause localization tool based on frequency-domain attention mechanisms, which identifies causal dependency patterns in frequency-domain features while filtering out low-frequency noise, achieving precise root cause localization.
  • Experiments on two benchmark datasets demonstrate that TAMO outperforms all compared methods including state-of-the-art (SOTA) approaches and derived baselines. Ablation and case studies further validate the effectiveness of the framework.
@@ II Related Work
In modern cloud native systems, with the widespread adoption of microservice architectures and dynamic infrastructures, it has become particularly important to ensure high system availability and rapid fault recovery, making root cause analysis (RCA) a critical part of operations and maintenance. As a result, various RCA methods have been developed to guarantee system reliability, including traditional RCA methods, multi-modal RCA methods, and LLM-based RCA methods.
@@
Traditional RCA methods: Traditional RCA methods are mainly based on statistical correlation analysis techniques, such as anomaly detection based on predefined rules[37] and root cause localization strategies based on causal discovery[7]. While these approaches perform well in static environments, they have difficulties in adapting to the dynamic service dependencies that characterize cloud native environments. With the rapid advancement of deep learning, neural network-based methods for real-time system data analysis continue to emerge. DeepLog[11] models log features via long short-term memory networks; GDN[10] and GTA[8] construct graph neural networks using metrics data to capture fault propagation paths; Sage[13] leverages causal Bayesian networks with trace data for root cause localization. However, these approaches predominantly focus on unimodal data, failing to exploit complementary inter-modal relationships.
@@
Multi-modal RCA methods: The multi-modal-based approach significantly reduces the missed detection rate by comprehensively considering system observation data from different modalities. Due to the differences between modalities, it is crucial to extract the correlated features across different data. PDiagnose[18] analyzes different modality data as independent entities and uses a voting mechanism to identify the root cause entity. Although this method successfully utilizes data from multiple modalities, it overlooks the consistency relationships between features from different modalities. Eadro[23], performs gated fusion of features extracted from different modalities and uses a graph neural network to fuse embedded features at the service level for root cause localization. HolisticRCA[14], compared to Eadro, fully considers the heterogeneous nature of microservice systems and standardizes the embedding of different modality data using an assembling building blocks strategy, combining masked embeddings for overall root cause analysis. In contrast to directly using raw data for root cause localization, Nezha[40] and DiagFusion[42] convert heterogeneous multimodal data into homogeneous event representations, performing joint analysis through event graphs to achieve root cause localization. CoE[38] further supports the operational experience of site reliability engineers as input on the event graph, enhancing the understanding and interpretability of the event. While these methods significantly improve the accuracy of root cause localization through multimodal fusion and comprehensive system analysis, their root cause analysis process still relies on human intervention by engineers.
LLM-based RCA methods: Due to the advanced reasoning and contextual understanding capabilities of large language models, LLM-based RCA methods have become significant advancements in automated root cause analysis. Currently, LLM-based RCA methods are still in the early stages of development. RCAgent[35] handles anomalies through a tool-augmented agent framework, while mABC[43] employs a blockchain-based multi-agent system for root cause voting and solution formulation. However, directly applying LLMs to RCA presents several critical challenges: the limited context window of LLMs restricts the processing of large-scale observation data, and due to text input constraints, they are unable to effectively handle time-series data (e.g., metrics), resulting in the loss of important feature information. Additionally, another key issue is how to use large models to capture and model dynamic dependencies in cloud native systems in real-time, without introducing misleading or false conclusions. These challenges highlight the need for further innovations in integrating LLM with multimodal and heterogeneous data sources for more accurate root cause analysis.
@@
Figure 2: The proposed TAMO framework consists of a Data Feature Observation Tool ($\mathscr{T}_{1}$), a Root Cause Localization Tool ($\mathscr{T}_{2}$), a Fault Classification Tool ($\mathscr{T}_{3}$), and an RCA Expert Agent ($\mathscr{A}$). In the framework, the agent $\mathscr{A}$ calls tools $\mathscr{T}_{1}$-$\mathscr{T}_{3}$ as perception tools to analyze the system contextual observational data in real time. The results of these perceptions are structured into text inputs that are fed back into the agent. By analyzing the perception results, the expert agent will provide the corresponding root cause analysis and repair suggestions.
@@ III METHODOLOGY
To fully exploit the multimodal observability data of cloud native systems while integrating prior domain knowledge, we propose TAMO, Tool-Assisted LLM Agent that Integrates Multi-Modality Observation Data to address root cause entity localization and fault classification in large-scale system operations and maintenance. This framework utilizes observation tools composed of a dual-branch collaborative diffusion model to effectively extract and integrate multimodal data from different entities. Additionally, task-specialized tools explicitly model the complex dependencies among entities using a causal graph, with feature propagation through the causal graph enabling accurate root cause analysis. Furthermore, to fully incorporate domain knowledge, we employ a LLM with operational expertise as an expert agent. The expert agent $\mathscr{A}$ integrates the outputs from the preceding tools $\mathscr{T}_{2}$-$\mathscr{T}_{3}$ with context knowledge of the system to offer human engineers a comprehensive fault analysis and recommended solutions. As shown in Fig.2, the proposed framework consists of three tools and one agent, $\mathscr{T}_{1}$$\mathscr{T}_{3}$ and $\mathscr{A}$, each responsible for a specific task in the root cause analysis process:
@@
    Data Feature Observation Tool ($\mathscr{T}_{1}$): This tool employs a dual-branch collaborative diffusion model to integrate multimodal features. It reconstructs and fuses textual data features, such as logs, into time-series data based on conditional guidance;
@@
    Fault Classification Tool ($\mathscr{T}_{3}$): Based on the causal relationships between entities and the invocation chain structure, this tool classifies the types of faults;
@@
    RCA Expert Agent ($\mathscr{A}$): Leveraging the expertise of a LLM, this agent synthesizes the outputs from previous tools along with system environment context to provide site reliability engineers (SREs) with a comprehensive fault analysis and recommended solutions.
III-A Data Feature Observation Tool ($\mathscr{T}_{1}$)

### v4

@@ Abstract
Implementing large language models (LLMs)-driven root cause analysis (RCA) in cloud-native systems has become a key topic of modern software operations and maintenance. However, existing LLM-based approaches face three key challenges: multi-modality input constraint, context window limitation, and dynamic dependence graph. To address these issues, we propose a tool-assisted LLM agent with multi-modality observation data for fine-grained RCA, namely TAMO, including multi-modality alignment tool, root cause localization tool, and fault types classification tool. In detail, TAMO unifies multi-modal observation data into time-aligned representations for cross-modal feature consistency. Based on the unified representations, TAMO then invokes its specialized root cause localization tool and fault types classification tool for further identifying root cause and fault type underlying system context. This approach overcomes the limitations of LLMs in processing real-time raw observational data and dynamic service dependencies, guiding the model to generate repair strategies that align with system context through structured prompt design. Experiments on two benchmark datasets demonstrate that TAMO outperforms state-of-the-art (SOTA) approaches with comparable performance.
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
@@ I Introduction
In recent years, microservice architecture in cloud-native systems has become foundational elements in modern enterprise software development, driving the rapid improvement of distributed systems [25][30][37][45]. Microservices decompose traditional monolithic applications into independent distributed services, while cloud native technologies leverage containerization and container orchestration to enhance the capabilities for flexible resource allocation and autoscaling [21][33]. In detail, cloud-native systems are composed of numerous independent microservices and containerized entities (e.g., pod, node)[52], which inevitably brings inter-service dependencies and complex communication patterns. Therefore, the failure of any entity (e.g., network corruption) can trigger cascading effects that damage overall system stability, potentially leading to significant economic losses[2][38].
@@
Root cause analysis (RCA) aims to quickly correlate features of multiple related entities, localize root cause, identify corresponding fault type, and provide reasonable solutions, thus realizing an automatic fault response mechanism and minimizing human intervention [5][48]. Therefore, how to perform accurate and efficient root cause analysis in cloud-native systems has become an important research topic recently [14][51]. On one hand, traditional RCA methods [29] [31] typically rely on rule-based anomaly detection or statistical correlation analysis (e.g. causal graphs)[24] to localize faults. These methods struggle to jointly analyze heterogeneous data modalities with unstructured logs and metric data, often leading to undetected faluts. On the other hand, deep learning (DL) based RCA methods attempt to incorporate multimodal feature fusion and dependency-sensitive graph neural networks (GNNs)[23][50] to achieve accurate fault localization. For instance, MULAN [51] infers causal structures from heterogeneous observation data for service-level diagnosis, while PDiagnose [18] treats multimodal data independently and aggregates decisions via voting to pinpoint root causes. The abovementioned methods mainly focus on specific tasks such as fault localization in complex systems, ignoring fine-grained RCA with fault type identification, contextual anomaly interpretation, and cross-level dependency reasoning. Recently, large language models (LLMs) have demonstrated revolutionary capabilities in automated RCA, opening new avenues for AIOps [9][12]. LLMs can combine RCA with current system context to generate reasonable fault diagnosis reports, emerging with RCACopilot[6]. Recent works such as RCAgent[41] and mABC[49] develop LLM-driven automated RCA frameworks that enable accurate fault localization and provide reasonable repair suggestions, thus further advancing the field of automated operations and maintenance. RCAgent proposes a tool-augmented autonomous agent framework that leverages log and code data with an internally deployed LLM to perform root cause analysis, while mABC mitigates circular dependency issue in microservice RCA through blockchain-inspired multi-agent voting within a structured workflow.
However, existing LLM-based methods still face some tricky challenges when applied to fine-grained RCA in cloud-native systems involving multi-level entities and heterogeneous multi-modality data:
  • Multi-Modality Input: Existing LLM-driven methods such as COCA[26], RCAgent[41], only support log data or code data as input, ignoring abundant context information such as metric data and trace data. However, directly feeding multimodal data into LLMs often hinders accurate RCA due to inherent modality misalignment, for instance, the semantic gap between time-series metrics and textual logs. Therefore, how to map different modalities into unified semantic space to assist the LLM in accurate RCA is challenging.
  • Context Window Limitation: Limited by fixed-length context windows, the LLM is unable to effectively process the massive real-time metric data generated by continuous monitoring across kinds of microservices, nodes or pods in cloud-native systems. The real-time metric data typically reflect system states but does not explicitly indicate anomalies or fault types, yet lacking aligned semantic meaning with text logs. So how to transform raw measurements into meaningful diagnostic prompts and input them into the LLM with limited context length?
  • Dynamic Dependence Graph: The dynamic service call chains from the trace data in the system inevitably bring evolving dependencies. However, exsiting LLM-based methods [40] struggle with fine-grained RCA over dynamic dependence graphs, making it difficult to capture implicit fault propagation path. So how to assist the LLM agent in modeling fault propagation paths based on dynamic dependence graph to improve root cause localization and fault types classification?
@@
To address these chanllenges, we propose a tool-assisted LLM agent framework based on kinds of domain-specific tools. This framework decouples the LLM from the raw multi-modality observation data by utilizing specialized tools: multi-modality alignment tool $\mathscr{T}_{1}$, root cause localization tool $\mathscr{T}_{2}$, and fault types classification tool $\mathscr{T}_{3}$. These specialized model tools are used to perceive the contextual environment, and their results form structured prompts that are fed back to the RCA expert agent for further analysis, generating accurate root cause analysis results and reasonable maintenance recommendations. In detail, multi-modality alignment tool $\mathscr{T}_{1}$ is based on a dual-branch diffusion architecture, which unifies multi-modality observation data, such as log data, trace data and metric data, into temporally consistent representations. To address dynamic dependencies challenge, the root cause localization tool ($\mathscr{T}_{2}$) jointly utilizes the extracted representations from $\mathscr{T}_{1}$ and dependence graphs extracted from trace data to model fault propagation and identify the root cause via frequency-domain attention-driven temporal causal analysis. Subsequently, the fault types classification tool ($\mathscr{T}_{3}$) further utilizes the casual graph from $\mathscr{T}_{2}$ to identify the anomalous patterns.
The main contributions are summarized as follows:
  • We propose a domain-specific tools assisted LLM agent framework for fine-grained RCA, including multi-modality alignment tool, root cause localization tool, and fault types classification tool to empower the LLM for real-time system diagnosis.
@@
  • In detail, we present a dual-branch collaborative diffusion tool for multi-modality alignment that captures potentially consistent patterns across modalities and provides unified representations for downstream tasks through condition-oriented collaborative reconstruction. Otherwise, the root cause localization tool is based on frequency-domain attention mechanisms, which identifies causal dependency patterns in frequency-domain features while filtering out low-frequency noise, achieving precise root cause localization.
  • Experiments on two benchmark datasets demonstrate that our proposed TAMO outperforms state-of-the-art (SOTA) approaches, achieving average Acc@1 improvement of $4.8\%$ for root cause localization and MiPr improvement of $10.8\%$ for fault types classification respectively.
@@ II Related Work
In modern cloud-native systems, with the widespread adoption of microservice architectures and dynamic infrastructures, it has become particularly important to ensure high system availability and rapid fault recovery, making root cause analysis (RCA) a critical part of operations and maintenance. As a result, various RCA methods have been developed to guarantee system reliability, including traditional RCA methods, multi-modal RCA methods, and LLM-based RCA methods.
@@
Traditional RCA methods: Traditional RCA methods are mainly based on statistical correlation analysis techniques, such as anomaly detection based on predefined rules and root cause localization strategies based on causal discovery[7][43]. While these approaches perform well in static environments, they have difficulties in adapting to the dynamic service dependencies that characterize cloud native environments. With the rapid advancement of deep learning, neural network-based methods for real-time system data analysis continue to emerge. DeepLog[11] models log features via long short-term memory networks. GDN[10] and GTA[8] construct graph neural networks using metrics data to capture fault propagation paths. Sage[13] leverages causal Bayesian networks with trace data for root cause localization. However, these approaches predominantly focus on unimodal data, failing to exploit complementary inter-modal relationships.
@@
Multi-modal RCA methods: The multimodal-based approach significantly reduces the undetected fault rate by comprehensively considering system observation data from different modalities. MULAN [51] learns causal structures from logs and metrics for service-level localization. PDiagnose[18] analyzes different modality data as independent entities and uses a voting mechanism to identify the root cause entity. Although this method successfully utilizes data from multiple modalities, it overlooks the consistency relationships between features from different modalities. Eadro[23] performs gated fusion of features extracted from different modalities and uses a graph neural network to fuse embedded features at the service level for root cause localization. HolisticRCA[14] compared to Eadro, fully considers the heterogeneous characteristics of cloud native systems and standardizes the embedding of different modality data using an assembling building blocks strategy, combining masked embeddings for overall root cause analysis. In contrast to directly using raw data for root cause localization, Nezha[46] and DiagFusion[48] convert heterogeneous multimodal data into homogeneous event representations, performing joint analysis through event graphs to achieve root cause localization. CoE[44] further supports the operational experience of site reliability engineers as input on the event graph, enhancing the understanding and interpretability of the event.
LLM-based RCA methods: Large language models are increasingly leveraged for root cause analysis due to their strong reasoning capabilities. Recent works explore various LLM applications: Adarma [36] combines LLMs with traditional ML for microservice anomaly detection and remediation using logs and metrics. RCAcopilot [6] maps multimodal observation data to predefined events and predicts root cause categories in production through an LLM-driven workflow, and agent-based approaches [35] utilize tool-assisted LLMs to actively retrieve diagnostic data like logs and metrics for localization. Despite these advancements, significant challenges persist. The limited context window of LLMs restricts their ability to process large scale of observation data, while their inherent text-based input modality makes handling time series metrics difficult, risking loss of critical temporal patterns. Furthermore, existing methods often face limitations in generalizability (e.g., reliance on predefined rules [6]),lack integration of crucial trace data for precise localization [35][36]. These challenges highlight the need for frameworks that can effectively unify multimodal data (logs, metrics, traces) within an LLM architecture to achieve both fine-grained root cause localization and accurate classification. We compare these existing methods with our approach in Table I.
@@
Figure 2: The proposed TAMO framework consists of a Multi-modality Alignment Tool ($\mathscr{T}_{1}$), a Root Cause Localization Tool ($\mathscr{T}_{2}$), a Fault Types Classification Tool ($\mathscr{T}_{3}$), and an RCA Expert Agent ($\mathscr{A}$). In the framework, the agent $\mathscr{A}$ calls tools $\mathscr{T}_{1}$-$\mathscr{T}_{3}$ as perception tools to analyze the system contextual observation data in real time. The results of these perceptions are structured into text inputs that are fed back into the agent. By analyzing the perception results, the expert agent will provide the corresponding root cause analysis and repair suggestions.
@@ III METHODOLOGY
TABLE I: Comparison of Different LLM RCA Models by Modality and Task. FTC denotes Fault Types Classification; RCL denotes Root Cause Localization
Methods Modality Task
Logs Metrics Traces FTC RCL
RCAcopilot[6] $\times$
Agent Work[35] $\times$ $\times$
Adarma[36] $\times$ $\times$ $\times$
Ours
To fully exploit the multi-modality observation data of cloud native systems while integrating prior domain knowledge, we propose TAMO, Tool-Assisted LLM Agent that Integrates Multi-Modality Observation Data to address root cause entity localization and fault types classification in cloud-native systems. This framework utilizes observation tools composed of a dual-branch collaborative diffusion model to effectively extract and integrate multimodal data from different entities. Additionally, task-specialized tools explicitly model the complex dependencies among entities using a causal graph, with feature propagation through the causal graph enabling accurate root cause analysis. Furthermore, to fully incorporate domain knowledge, we employ a LLM with operational expertise as an expert agent. The expert agent $\mathscr{A}$ integrates the outputs from the preceding tools $\mathscr{T}_{2}$-$\mathscr{T}_{3}$ with context knowledge of the system to offer human engineers a comprehensive fault analysis and recommended solutions. As shown in Fig.2, the proposed framework consists of three tools and one agent, $\mathscr{T}_{1}$$\mathscr{T}_{3}$ and $\mathscr{A}$, each responsible for a specific task in the root cause analysis process:
@@
    Multi-modality Alignment Tool ($\mathscr{T}_{1}$): This tool employs a dual-branch collaborative diffusion model to integrate multimodal features. It reconstructs and fuses textual data features, such as logs, into time-series data based on conditional guidance;
@@
    Fault Types Classification Tool ($\mathscr{T}_{3}$): Based on the causal relationships between entities and the calling chain topology, this tool classifies the types of faults;
@@
    RCA Expert Agent ($\mathscr{A}$): Leveraging the expertise of a LLM, this agent synthesizes the outputs from previous tools along with system environment context to provide a comprehensive fault analysis and recommended solutions.
III-A Multi-modality Alignment Tool ($\mathscr{T}_{1}$)